Three consecutive integers add up to 57. What are these integers?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us say the three consecutive numbers be x-1, x and x+1

Now, as per the given question,

x-1+x+x+1=57

3x=57

x=57/3

x=19

Hence,

x-1=19-1=18

x=19

x+1=19+1=20

Thus , the three consecutive numbers are 18,19 and 20.

Thanks